Sage Company reported the following amounts in the stockholders’ equity section of its December 31, 2016, balance sheet.

Preferred stock, 10%, $100 par (10,000 shares authorized, 1,800 shares issued) $180,000
Common stock, $5 par (103,500 shares authorized, 20,700 shares issued) 103,500
Additional paid-in capital 114,000
Retained earnings 423,000
   Total $820,500


During 2017, Sage took part in the following transactions concerning stockholders’ equity.

1. Paid the annual 2016 $10 per share dividend on preferred stock and a $2 per share dividend on common stock. These dividends had been declared on December 31, 2016.
2. Purchased 1,500 shares of its own outstanding common stock for $43 per share. Sage uses the cost method.
3. Reissued 600 treasury shares for land valued at $27,100.
4. Issued 550 shares of preferred stock at $105 per share.
5. Declared a 10% stock dividend on the outstanding common stock when the stock is selling for $41 per share.
6. Issued the stock dividend.
7. Declared the annual 2017 $10 per share dividend on preferred stock and the $2 per share dividend on common stock. These dividends are payable in 2018.

Please prepare the December 31, 2017, stockholders’ equity section. Assume 2017 net income was $336,000.

SAGE COMPANY
Stockholders’ Equity

For the Year Ended December 31, 2017

Capital Stock

Preferred Stock $235,000

Common Stock $?

Total Capital Stock $?

Additional Paid-in Capital $?

Total Paid-in Capital $?

Retained Earnings $?

Total Paid-in Capital Retained Earnings $?

Less: Treasury Stock $38,700

Total Stockholder's Equity $?
Answer #1
SAGE COMPANY
Stockholders’ Equity
For the Year Ended December 31, 2017
Capital Stock
Preferred Stock (180000+550*100) $ 235000
Common Stock (9900+103500) 113400
Total Capital Stock 348400
Additional Paid-in Capital (114000+1300+550*5+1980*36) 189330
Total Paid-in Capital 537730
Retained Earnings   610760
Total Paid-in Capital Retained Earnings 1148490
Less: Treasury Stock (1500-600)*43 (38700)
Total Stockholder's Equity $ 1109790

Stock dividend = (20700-1500+600)*10%= 1980

Retained earnings= 423000+336000-1980*41-(550+1800)*10+(19800+1980)*2=